DETAILS LARLINALG8 8.4.039. Find the Euclidean distance between u and v. - (6,0), v = (1,1) d(u, v) =
Answer #1

U=(6,0), V=(1,1) du,v) = 5(6-7) (0-1) = 36-121 + i tri [:12-1] = 36-121-2 534-121